# DreamDrone

### v2

@@ 3.3 Wrapping latent codes
Algorithm 1 Low-pass latent code wrapping
$\boldsymbol{x}_{t}$$\triangleright$ latent code at timestep $t$ of current view $c$
1:$F(\boldsymbol{x}_{t})\leftarrow FFT(\boldsymbol{x}_{t})$triangleright Apply Fast Fourier Transform
2:Split $F(\boldsymbol{x}_{t})$ into $F_{low}$ and $F_{high}$ using threshold $\sigma$
3:$\boldsymbol{x}_{t}^{low}\leftarrow IFFT(F_{low})$triangleright Inverse FFT on low-frequency component
4:$\boldsymbol{x}_{t}^{low-wrapped}\leftarrow Wrap(\boldsymbol{x}_{t}^{low})$triangleright Wrap the low-frequency content
5:$F_{wrapped}\leftarrow FFT(\boldsymbol{x}_{t}^{low-wrapped})$triangleright FFT on wrapped content
@@
6:$F^{\prime}\leftarrow F_{wrapped}+F_{high}$triangleright Combine low-frequency of wrapped content with high-frequency of original code
[…]
@@
The results in the right side of Fig. 2 reveal that directly wrapping images based on camera intrinsics $K$, extrinsics ($R,t$), and depth information leads to regions of distortion in the images. Additionally, the use of inpainting [34, 22, 48] and outpainting [16, 46] models to fill these gaps does not achieve satisfactory outcomes. In pursuit of generating photo-realistic images, and considering the robustness of the denoising process and decoder in the stable diffusion model, we opt to edit the latent code corresponding to timestep $t$. PnP [39] and DIFT [38] have shown that the features of diffusion possess strong semantic information, with semantic parts being shared across images at each step. The simplest method for wrapping the latent code follows the same approach as wrapping the image. The only difference between wrapping the latent code and wrapping the image is a slight modification in the camera intrinsics; this entails scaling the camera intrinsics proportionally based on the different resolutions of the image and latent code.
@@
The overall procedure for wrapping the latent code is illustrated in Alg. 1. Initially, a latent code ${x}_{t}$ is obtained and transformed via Fast Fourier Transform (FFT) to $F(\boldsymbol{x}_{t})$. This is divided into low-frequency $F_{\text{low}}$ and high-frequency $F_{\text{high}}$ components, segregated at threshold $\lambda$. The key step involves wrapping the Inverse FFT (IFFT) processed low-frequency component $x_{t}^{\text{low}}=\text{IFFT}(F_{\text{low}})$, forming $\boldsymbol{x}_{t}^{low-wrapped}$. Merging $\text{FFT}(\boldsymbol{x}_{t}^{low-wrapped})$ with $F_{\text{high}}$, we obtain $F_{\text{new}}$, from which the final latent code $\boldsymbol{x}_{t}^{\prime}=\text{IFFT}(F^{\prime})$ is reconstructed. This approach efficiently preserves high-frequency details, enabling high-fidelity scene generation aligned with text prompts.
@@ 3.4 Feature-correspondence-guidance design
To maintain consistency between the generated result and the original image, inspired by recent image and video editing works [39, 45, 8, 42], we modify the process of the self-attention module of U-Net when denoising the latent code $\boldsymbol{x}_{t}^{\prime}$. Specifically, we denoise the views for the current and next view together. The key and value of the self-attention modules from the next view are replaced by that of the current view. To be specific, for obtaining the original view, the self-attention module is defined the same as Eq. 4. The modified cross-view self-attention for generating a novel view is defined as:
$\boldsymbol{o}^{\prime}=\mathbf{A^{\prime}}\mathbf{V},\;\;\textup{where}\;\mathbf{A}^{\prime}=\textup{Softmax}(\mathbf{Q}^{\prime}\mathbf{K}^{\top}),$ (5)
where $\mathbf{Q}^{\prime}$, $\mathbf{A}^{\prime}$, and $\boldsymbol{o}^{\prime}$ are query, attention matrix, and output features for the novel views. $\mathbf{K}$ and $\mathbf{V}$ are injected keys and values obtained from the self-attention module for generating the original view.
@@
Maintaining geometry consistency between adjacent views using the cross-view self-attention mechanism presents challenges, especially in preserving high-frequency details as the camera moves forward. The recent DIFT study [38] highlights the potential of using intermediate features of diffusion models for accurate point-to-point image matching. Additionally, the concept of vanilla classifier guidance [5] steers the diffusion sampling process using pre-trained classifier gradients towards specific class labels. Building on these ideas, we integrate feature correspondence guidance into the DDIM process to enhance consistency between adjacent views, addressing the challenge of detail preservation in dynamic scenes.
Specifically, we obtain the features of the current and next view at each timestep $t$ of the DDIM process and calculate the cosine distance between the wrapped original features and features from the next novel views:
$\mathcal{L}_{sim}^{t}=\frac{1-\cos\left[\mathrm{Wrap}(f_{t}),f_{t}^{\prime}\right]}{2},$ (6)
where $f_{t}$ and $f_{t}^{\prime}$ are diffusion features extracted from pre-trained U-Net $\boldsymbol{\epsilon}_{\theta}$ at timestep $t$ and $\mathrm{Wrap}$ is the wrapping functions. The lower $\mathcal{L}_{sim}^{t}$, the higher the similarity.
We further introduce the similarity score $\mathcal{L}_{sim}$ to the DDIM sampling process, for generating novel views with geometry consistency. The predicted noise $\hat{\epsilon}$ is formulated as:
$\hat{\epsilon}=\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t})-\lambda\sqrt{\bar{\alpha}_{t-1}}\triangledown_{\boldsymbol{x}_{t}}\mathcal{L}_{sim}^{t},$ (7)
where $\lambda$ is the constant hyper-parameter and latent code $\boldsymbol{x}_{t-1}$ is calculated by Eq. 3
@@ 4 Experiments
Figure 3: Qualitative comparisons of InfNat-0 [17] and ours. We provide four scene images with various styles and categories as start points and ask models to fly through the images. 10 frames are illustrated for each starting point.
Implementation details.
We take Stable Diffusion [34] with the pre-trained weights from version 2.1https://huggingface.co/stabilityai/stable-diffusion-2-1-base as the basic text-to-image diffusion and MiDas [32] with weights $\mathrm{dpt\_beit\_large\_512}$https://github.com/isl-org/MiDaS. The overall diffusion timesteps is $1000$. We wrap the latent code at timestep $t_{1}$=$21$ and add more degrees of noise to timestep $t_{2}$=$441$. The threshold $\sigma$ for low-pass wrapping is $20$ and the hyper-parameter $\lambda$ for feature-correspondence guidance is $300$. Due to the page limit, please refer to the supplementary material (supp.) for details.
Baselines.
@@
We compare against 1) 2 supervised methods for perpetual view generation: InfNat [19] and InfNat-0 [17]. 2) one text-conditioned scene generation: SceneScape [7]. 3) 2 supervised methods for text-to-video generation: CogVideo [11] and VideoFusion [23]. 4) one method for zero-shot text-to-video generation: T2V-0 [14].
Evaluation Metrics.
@@
The perpetual view generation is evaluated both quantitatively and qualitatively. We evaluate CLIP [31] score, which indicates text-scene alignment for quantitative comparisons. CogVideo [11], VideoFusion [23], SceneScape [7], and T2V-0 [14] are all engaged in text-conditioned generation tasks. We generated 50 scene-related text prompts using GPT-4https://openai.com/gpt-4 and then created videos using each of the three methods. For the InfNat [19] and InfNat-0 [17] methods, we used Stable Diffusion to generate the initial frame, followed by subsequent frame generation based on this initial frame. We calculated the distance between each generated frame and the text embedding, known as the CLIP score. A high average CLIP score indicates not only that the generated images are more aligned with the corresponding prompts but also that they consistently maintain high quality. Considering that the InfNat [19] and InfNat-0 [17] methods might gradually generate other scenes, we provided 10 very general prompts such as ‘an image of the landscape’ and ‘an image of the mountain’ for these methods, and then selected the highest CLIP score as the CLIP score for the current frame.
@@ 4.1 Quantitative comparison
The quantitative comparison results are shown in Tab. 1. It is evident that the CLIP scores for the InfNat and InfNat-0 methods are low, indicating poor robustness. These methods are not sufficiently versatile to handle infinite scene generation for any scenario, as also illustrated by the visual results in Fig. 3. CogVideo and VideoFusion, which are supervised text-to-video generation methods, struggle to generate first-person perspective landscapes due to limitations of text-video training datasets. Additionally, their CLIP scores decrease with an increasing number of frames, such as 32 frames, which is corroborated by the visual results in supp. T2V-0 [14], SceneScape [7], and our method, both based on the pre-trained stable diffusion model and training-free, have relatively higher CLIP scores. However, as the number of frames increases, SceneScape and T2V-0’s CLIP scores decrease, whereas ours remains constant (please refer to Fig. 4 for qualitative comparisons). We attribute this to our low-pass wrapping strategy, which provides better content guidance for the generated images, and the introduction of feature-correspondence guidance, ensuring the coherence of adjacent frames and thus maintaining a higher CLIP score.
Table 1: Quantitative comparison results. We calculate the average CLIP score [31] across various video lengths.
8 frames 16 frames 32 frames
InfNat [19] 0.125 0.123 0.118
InfNat-0 [17] 0.128 0.125 0.122
SceneScape [7] 0.263 0.238 0.174
CogVideo [11] 0.255 0.249 0.241
VideoFusion [23] 0.281 0.283 0.272
T2V-0 [14] 0.312 0.305 0.287
Ours 0.320 0.318 0.319
4.2 Qualitative comparison
T2V-0 [14] introduces unsupervised text-conditioned video generation using stable diffusion. SceneScape [7] focuses on ‘zoom out’ effects during backward camera movement. However, as seen in Fig. 4, both methods have limitations. SceneScape struggles with outdoor scenes and forward camera movement, leading to blurred and distorted results after 8 steps due to its reliance on a pre-trained inpainting model. T2V-0 displays a decline in quality beyond the third frame in complex environments like Lego-style cities, likely from its latent code editing approach that compromises frame continuity and geometric consistency. Conversely, our DreamDrone excels across various scenes. It maintains detail, continuity, and quality in advancing camera scenarios, evident in even simpler landscapes like mountains where T2V-0 and SceneScape cannot effectively portray dynamic elements like cloud movement. Our approach ensures the preservation of fine details such as shadows and sunlight, creating a more dynamic and realistic video experience. Please refer to supp. for more quantitative comparisons.
Our task bears similarities to text-to-video generation, we further provide qualitative comparisons with VideoFusion [23], one of the state-of-the-art methods for text-to-video generation. Due to the page limit, please refer to supp. for detailed comparisons.
4.3 Ablation studies
Figure 5: Ablation study on the key components of our method. We perform ablation studies by disabling the key components of our method: feature-correspondence guidance, low-pass wrapping strategy, and cross-view self-attention module.
We perform ablation studies on 1) low-pass feature wrapping, 2) cross-view self-attention module, and 3) feature-correspondence guidance in Fig. 5. Without feature-correspondence guidance and cross-view attention, even if we perform the corresponding wrapping at the feature level, it is challenging to ensure strong consistency between generated adjacent frames ($1^{st}$ row). Direct wrapping of the latent code leads to increasingly blurred images and a loss of high-frequency details. However, the use of our low-pass wrapping strategy effectively ensures that the generated images retain high-definition details ($2^{nd}$ row). Without the guidance of the feature-correspondence guidance mechanism or the presence of a cross-view self-attention module, it is difficult to guarantee the geometric consistency of adjacent frames ($3^{rd}$ and $4^{th}$ rows). The components in our proposed pipeline enhance each other, ensuring that each generated frame is not only rich in detail but also maintains good consistency with adjacent frames. Please refer to supp. for more detailed ablation studies.

### v3

@@ 3.3 Warping latent codes
Algorithm 1 Warping latent code with high-pass filter
1:$\bm{x}_{t}$triangleright latent code at timestep $t$ of current view $c$
2:$F(\bm{x}_{t})\leftarrow FFT(\bm{x}_{t})$triangleright Apply Fast Fourier Transform
3:Split $F(\bm{x}_{t})$ into $F_{low}$ and $F_{high}$ using threshold $\sigma$
4:$\bm{x}_{t}^{low}\leftarrow IFFT(F_{low})$triangleright Inverse FFT on low-frequency component
5:$\bm{x}_{t}^{low-warped}\leftarrow warp(\bm{x}_{t}^{low})$triangleright warp the low-frequency content
6:$F_{warped}\leftarrow FFT(\bm{x}_{t}^{low-warped})$triangleright FFT on warped content
@@
7:$F^{\prime}\leftarrow F_{warped}+F_{high}$triangleright Combine low-frequency of warped content with high-frequency of original content
8:$\bm{x}_{t}^{\prime}\leftarrow IFFT(F^{\prime})$triangleright Inverse FFT to get latent code for next view $c^{\prime}$ return $\bm{x}_{t}^{\prime}$triangleright warped latent code at timestep $t$ for next view $c^{\prime}$
@@
The results in the right side of Fig. 2 reveal that directly warping images based on camera intrinsics $K$, extrinsics $\left\{\mathcal{R},\mathcal{T}\right\}$, and depth information leads to regions of distortion in the images. Additionally, the use of inpainting [36, 23, 53] and outpainting [17, 49, 51] models to fill these gaps does not achieve satisfactory outcomes. In pursuit of photo-realistic images, we opt to edit the latent code corresponding to timestep $t$. PnP [44] and DIFT [42] have shown that the features of diffusion possess strong semantic information, with semantic parts being shared across images at each step. The simplest method for warping the latent code follows the same approach as warping the image. The only difference between warping the latent code and warping the image is a slight modification in the camera intrinsics; this entails scaling the camera intrinsics proportionally based on the different resolutions of the image and latent code.
@@
The overall procedure for warping the latent code is illustrated in Alg. 1. Initially, a latent code ${x}_{t}$ is obtained and transformed via Fast Fourier Transform (FFT) to $F(\bm{x}_{t})$. This is divided into low-frequency $F_{\text{low}}$ and high-frequency $F_{\text{high}}$ components, segregated at threshold $\sigma$. The key step involves warping the Inverse FFT (IFFT) processed low-frequency component $\bm{x}_{t}^{\text{low}}=\text{IFFT}(F_{\text{low}})$, warping to the next view $\bm{x}_{t}^{low-warped}$. Merging $\text{FFT}(\bm{x}_{t}^{low-warped})$ with $F_{\text{high}}$, we obtain $F^{\prime}$, from which the final latent code $\bm{x}_{t}^{\prime}=\text{IFFT}(F^{\prime})$ is reconstructed. This approach efficiently preserves high-frequency details, enabling high-fidelity scene generation aligned with text prompts.
@@ 3.4 Feature-correspondence-guidance design
To maintain consistency between the generated result and the original image, inspired by recent image and video editing works [44, 48, 8, 45], we modify the process of the self-attention module of U-Net when denoising the latent code $\bm{x}_{t}^{\prime}$. Specifically, we denoise the views for the current and next view together. The key and value of the self-attention modules from the next view are replaced by that of the current view. To be specific, for obtaining the original view, the self-attention module is defined the same as Eq. 4. The modified cross-view self-attention for generating a novel view is defined as:
$\bm{o}^{\prime}=\mathbf{A^{\prime}}\mathbf{V},\;\;\textup{where}\;\mathbf{A}^{\prime}=\textup{Softmax}(\mathbf{Q}^{\prime}\mathbf{K}^{\top}),$ (5)
where $\mathbf{Q}^{\prime}$, $\mathbf{A}^{\prime}$, and $\bm{o}^{\prime}$ are query, attention matrix, and output features for the novel views. $\mathbf{K}$ and $\mathbf{V}$ are injected keys and values obtained from the self-attention module for generating the original view. Please note that the $\mathbf{K}$ and $\mathbf{V}$ are also warped before injection.
@@
Maintaining geometry consistency between adjacent views using the cross-view self-attention mechanism presents challenges, especially in preserving high-frequency details as the camera moves forward. The recent DIFT [42] highlights the potential of using intermediate features of diffusion models for accurate point-to-point image matching [32]. Additionally, the concept of vanilla classifier guidance [5] steers the diffusion sampling process using pre-trained classifier gradients towards specific class labels. Building on these ideas, we integrate feature correspondence guidance into the DDIM sampling process to enhance consistency between adjacent views, addressing the challenge of detail preservation in dynamic scenes.
Specifically, we obtain the features of the current and next view at each timestep $t$ of the DDIM process and calculate the cosine distance between the warped original features and features from the next novel views:
$\mathcal{L}_{sim}^{t}=\frac{1-\cos\left[\mathrm{warp}(f_{t}),f_{t}^{\prime}\right]}{2},$ (6)
where $f_{t}$ and $f_{t}^{\prime}$ are intermediate features extracted from pre-trained U-Net $\bm{\epsilon}_{\theta}$ at timestep $t$ and $\mathrm{warp}$ is the warping functions. The lower $\mathcal{L}_{sim}^{t}$, the higher the similarity.
We further introduce the similarity score $\mathcal{L}_{sim}^{t}$ to the DDIM sampling process, for generating novel views with geometry consistency. The predicted noise $\hat{\epsilon}$ is formulated as:
$\hat{\epsilon}=\bm{\epsilon}_{\theta}(\bm{x}_{t})-\lambda\sqrt{\bar{\alpha}_{t-1}}\triangledown_{\bm{x}_{t}}\mathcal{L}_{sim}^{t},$ (7)
where $\lambda$ is the constant hyper-parameter and latent code $\bm{x}_{t-1}$ is calculated by Eq. 3
@@ 4 Experiments
4.1 Implementation details
[…]
4.2 Baselines
@@
We compare against 1) two supervised methods for perpetual view generation: InfNat [20] and InfNat-0 [18]. 2) one text-conditioned 3D point cloud-based scene generation: SceneScape [7]. 3) two supervised methods for text-to-video generation: CogVideo [11] and VideoFusion [24]. 4) one method for zero-shot text-to-video generation: T2V-0 [14].
4.3 Evaluation metrics
We evaluate our zero-shot perpetual scene generation into two aspects: 1) the quality of generated images and text-image alignment, and 2) the temporal consistency of generated image sequences.
Image quality and text-image alignment.
@@ 4.3 Evaluation metrics
We evaluate CLIP score [33], which indicates text-scene alignment for quantitative comparisons. A high average CLIP score indicates not only that the generated images are more aligned with the corresponding prompts but also that they consistently maintain high quality [14]. CogVideo [11], VideoFusion [24], SceneScape [7], and T2V-0 [14] are all engaged in text-conditioned generation tasks. We generated 50 scene-related text prompts using GPT-4333https://openai.com/gpt-4 and then created videos using each of the three methods. For the InfNat [20] and InfNat-0 [18] methods, we used Stable Diffusion to generate the initial frame, followed by subsequent frame generation based on this initial frame. We calculated the distance between each generated frame and the text embedding, known as the CLIP score. Considering that the InfNat [20] and InfNat-0 [18] methods trained on natural scene datasets, we further provided 10 very general prompts such as ‘an image of the landscape’ and ‘an image of the mountain’ for these methods, and then selected the highest CLIP score as the CLIP score for the current frame.
Table 1: Ablations of image quality and temporal coherence of generated image sequences with various lengths. Please refer to Fig. 3 for quality comparisons.
Methods PSNR $\uparrow$ SSIM $\uparrow$ CLIP $\uparrow$
8 frames 16 frames 32 frames 8 frames 16 frames 32 frames 8 frames 16 frames 32 frames
warp image 26.90 22.46 21.62 0.25 0.23 0.24 0.138 0.112 0.106
warp latent 28.35 28.57 28.75 0.27 0.28 0.24 0.135 0.122 0.125
warp latent+DDPM 24.67 23.04 22.59 0.12 0.10 0.06 0.302 0.297 0.308
warp latent+DDPM+guidance 28.27 28.21 28.10 0.34 0.30 0.26 0.317 0.316 0.313
warp latent+DDPM+guidance +cross-view attn. 28.89 28.83 28.75 0.32 0.31 0.27 0.318 0.315 0.315
warp latent+DDPM+guidance +cross-view attn.+high pass filter 29.91 29.86 29.79 0.39 0.38 0.35 0.320 0.318 0.319
Temporal consistency of generated image sequences.
We demonstrate our advancements in temporal consistency against other SOTA methods by calculating average PSNR and SSIM scores across adjacent frames for generated videos with different lengths. The higher scores demonstrate the superiority in terms of cross-view consistency.
4.4 Ablation studies
We perform ablation studies on our three proposed modules: 1) warping latent with high-pass filter, 2) cross-view self-attention module, and 3) feature-correspondence guidance. The quantitative ablation results are shown in Tab. 1 and we visualize the ablation samples in Fig. 3.
Figure 3: Ablation results for the key components. We perform ablation studies by disabling the key components of our method. We illustrate every five frames for each ablation experiment. Please zoom in for better comparisons.
The simplest method for infinite scene generation tasks is frame-by-frame image warping, but this approach is unfeasible, as is directly warping the latent code. Warping images leads to non-integer pixel coordinates, resulting in interpolation-induced blurring and distortion. Moreover, these errors accumulate with each frame generated, leading to a collapse in quality. The first two rows of Table 1 show that directly warping images or warping latent codes (i.e., removing DDPM) results in very low CLIP scores, indicating poor quality of the generated images. The generated images becoming progressively blurred can also be observed in the first two rows of Fig. 3. We introduce DDPM to increase the degrees of freedom of the diffusion model, thereby generating high-quality images. However, the introduction of DDPM has the side effect of worsening the semantic consistency between adjacent frames ($3^{rd}$ row in Tab. 1 and Fig. 3).With the help of DDPM, the CLIP score increases from $0.125$ to $0.308$ when generating 32 images. Please refer to the supp. for the generated results with different scales of the DDPM forward process.
To ensure the quality of image generation while also maintaining consistency with adjacent views, we propose a feature-correspondence guidance strategy. Comparing the third and fourth rows of Fig. 3, it is evident that the semantic consistency between adjacent frames is significantly enhanced after adding guidance, with noticeable improvements in both PSNR and SSIM scores in Tab. 1. To further enhance cross-view consistency, we adopted cross-view attention modules and high-pass filtering. From the visualized results at the $5^{th}$ and $6^{th}$ rows in Fig. 3, it is clear that the semantic consistency of adjacent camera perspectives is further strengthened after incorporating the cross-view attention module. The operation of the high-pass filter further preserves the high-frequency details of the current frame, thereby further enhancing the semantic consistency of high-frequency details between adjacent frames. For instance, comparing the left side house at the $4^{th}$, $5^{th}$, and $6^{th}$ rows in Fig. 3, the cross-view consistency is enhanced after adding the proposed modules.
In addition to conducting ablation experiments on the modules we propose, we continue to explore two more questions:
Figure 4: Ablation study for scene travel. We visualize two image sequences and change the prompt when generating novel views. We illustrate every five images and the prompts are changed when generating $31^{th}$ image ($7^{th}$ image shown in each row).
Q1: Can DreamDrone shuttle from one scene to another by changing text prompts? During the frame-by-frame process, we changed the textual prompts, with the generated results shown in Fig. 4. The visualized results demonstrate that DreamDrone can smoothly complete the scene travel (from streets in Inazuma City to urban art street) or the transition of scene styles (from realistic to Lego style) while ensuring the semantic consistency of adjacent views, according to the changes in textual prompts.
Figure 5: Ablation study on customized camera trajectory. We generate images with different camera directions. For the sample of the Eiffel Tower, our camera perspective continuously ascends. For the $2^{nd}$ scene of the Lego city, our camera not only moves forward but also shifts upwards and to the right.
Q2: Can explicitly warping the latent code control the trajectory of camera perspective movement? Since our method generates image sequences frame by frame, we can freely adjust the camera’s flight angle by altering the camera’s extrinsic parameters. In Fig. 5, we provide sequences of images generated under different camera trajectories. The results show that our method possesses a high degree of freedom, allowing for the free customization of the camera’s trajectory. Other state-of-the-art methods cannot achieve this functionality.
4.5 Qualitative comparison
Figure 6: Qualitative comparisons of InfNat-0 [18] and ours. We provide four starting scene images with various styles and categories as start points and ask models to fly through the images. 50 frames are generated and we illustrate every five frames for each starting scene image.
T2V-0 [14] introduces unsupervised text-conditioned video generation using stable diffusion. SceneScape [7] focuses on ‘zoom out’ effects during backward camera movement. However, as seen in Fig. 7, both methods have limitations. SceneScape struggles with outdoor scenes and forward camera movement, leading to blurred and distorted results after 8 steps due to its reliance on a pre-trained inpainting model. T2V-0 displays a drop in quality beyond the third frame in complex environments like Lego-style cities, likely from its latent code editing approach that compromises frame continuity and geometric consistency. Conversely, our DreamDrone excels across various scenes. It maintains detail, continuity, and quality in advancing camera scenarios, evident in even simpler landscapes like mountains where T2V-0 and SceneScape cannot effectively portray dynamic elements like cloud movement. Our approach ensures the preservation of fine details such as shadows and sunlight, creating a more dynamic and realistic video experience. Please refer to supp. for more comparisons.
As our task bears similarities to text-to-video generation, we further provide qualitative comparisons with VideoFusion [24]. Due to the page limit, please refer to supp. for detailed comparisons.
@@ 4.6 Quantitative comparison
Tab. 2 offers a detailed comparison of various SOTA methods for generating image sequences, including our method, DreamDrone. When compared to other training-based methods, DreamDrone, despite being training-free, consistently achieves higher CLIP scores across all frame lengths (0.320, 0.318, 0.319 for 8, 16, and 32 frames respectively). This is particularly noteworthy as the CLIP scores for training-based methods generally degrade as the number of generated frames increases. For instance, VideoFusion’s [24] CLIP scores decrease from 0.281 for 8 frames to 0.272 for 32 frames. This trend suggests a decline in the quality of generated images with an increase in sequence length for training-based methods.
Table 2: Qualitative comparisons with other SOTA methods. We evaluate the quality and temporal coherence of the generated image sequences with various lengths.
Methods PSNR $\uparrow$ SSIM $\uparrow$ CLIP $\uparrow$
8 frames 16 frames 32 frames 8 frames 16 frames 32 frames 8 frames 16 frames 32 frames
training-based InfNat [20] 28.75 28.67 28.65 0.32 0.30 0.30 0.125 0.123 0.118
InfNat-0 [18] 28.92 28.89 28.87 0.37 0.35 0.34 0.128 0.125 0.122
CogVideo [11] 31.03 30.08 29.32 0.45 0.39 0.31 0.255 0.249 0.241
VideoFusion [24] 29.89 28.36 28.78 0.41 0.37 0.31 0.281 0.283 0.272
training-free T2V-0 [14] 27.25 26.17 26.03 0.27 0.24 0.23 0.312 0.305 0.287
Scenescape [7] 29.87 29.75 29.66 0.41 0.38 0.34 0.318 0.282 0.279
DreamDrone (Ours) 29.91 29.86 29.79 0.39 0.38 0.35 0.320 0.318 0.319
In contrast, DreamDrone maintains high CLIP scores even as the sequence length increases, indicating superior image quality. When compared to other training-free methods, DreamDrone also stands out. For example, while T2V-0’s [14] CLIP scores decrease from 0.312 for 8 frames to 0.287 for 32 frames, DreamDrone’s CLIP scores remain relatively stable, further demonstrating its robustness in maintaining image quality across varying sequence lengths. This analysis underscores the effectiveness of DreamDrone in generating high-quality, temporally coherent image sequences without the need for training.